OpenStudy (anonymous):

The radioactive isotope of lead, Pb-209, decays at a rate proportional to the amount present at time t and has a half life of 3.3 hours. if 1 gram of this isotope is present initially, how long will it take for 90% to decay

OpenStudy (anonymous):

\[\frac{dP}{dt}=kP\] given P(0)=1g P(3.3)=.5g P(x)=.1g

OpenStudy (anonymous):

doing the integration again you get \[P=ce^{kt}\]

OpenStudy (anonymous):

\[1=ce^0\] c=1

OpenStudy (anonymous):

\[P=e^{kt}\]

OpenStudy (cwrw238):

i think it should be dP/dt = -kP as this is decay

OpenStudy (anonymous):

does it matter as k is a constant

OpenStudy (anonymous):

i'm looking at an example now and the book uses kP

OpenStudy (cwrw238):

right - good point - its just that in all problems like this this is the way the textbooks do them

OpenStudy (cwrw238):

ooh - ok

OpenStudy (anonymous):

like i understand how to do it but i feel i messed up in simplifying because i get a negative time

OpenStudy (anonymous):

\[.5=e^{3.3k}\]

OpenStudy (anonymous):

\[ln(.5)=ln(e^{3.3k})\] \[ln(.5)=3.3k\] \[k=\frac{ln(.5)}{3.3}\]

OpenStudy (cwrw238):

try .5 = e ^(-3.3t)

OpenStudy (anonymous):

\[P=e^{\frac{ln(.5)t}{3.3}}\]

OpenStudy (cwrw238):

-3.3k i mean

OpenStudy (anonymous):

oh wait i found a simplifying error but i don't think it'll do anything

OpenStudy (anonymous):

\[P=e^{\frac{ln(.5)}{3.3}}e^t\]

OpenStudy (anonymous):

wups i meant this \[e^{\frac{ln(.5)}{3.3}}e^{\frac{t}{3.3}}\]

OpenStudy (anonymous):

\[{.5^{\frac{1}{3.3}}}e^\frac{t}{3.3}=p\]

OpenStudy (cwrw238):

i guess i'm pretty rusty at these....

OpenStudy (anonymous):

but lets see hwat i get with this

OpenStudy (anonymous):

\[\frac{.1}{.5^{\frac{1}{3.3}}}=e^{\frac{t}{3.3}}\]

OpenStudy (anonymous):

it's not even close to the answer lol

OpenStudy (anonymous):

@asnaseer

OpenStudy (asnaseer):

you have it correct up to this point:\[P=e^{\frac{ln(.5)t}{3.3}}\]you can then say the following:\[P=e^{\frac{ln(.5)t}{3.3}}=(e^{ln(.5)})^{\frac{t}{3.3}}=(0.5)^{\frac{t}{3.3}}\]

OpenStudy (asnaseer):

which can also be written as:\[P=\frac{1}{2^{\frac{t}{3.3}}}=2^{-\frac{t}{3.3}}\]

OpenStudy (anonymous):

still not gettin the right answer

OpenStudy (campbell_st):

ok... this is how to do it \[0.5 = 1\times e^{3.3k}\] solve for k... you'll need logs \[\ln(0.5) = 3.3k........ k = \frac{\ln(0.5)}{3.3} ..... k = -0.210045\] so your model is \[P = 1 \times e^{-0.210045t}\]

OpenStudy (campbell_st):

next find when P = 0.1 \[01. = 1\times e^{-0.210045t}\] take the log of both sides \[l(0.1) = -0.210045t....... t = \frac{\ln(0.1)}{-0.210045}.... t=10.9623\]

OpenStudy (campbell_st):

the time taken is 10.9623 hours for the isotope to decay to 0.1 grams from its original of 1 gram